# Temperature Dependent Mechanical and Structural Properties of Uniaxially Strained Planar Graphene

### v1

@@ 3.3 Bond length and bond angle distributions
In order to analyze the effects of temperature and stress on the distributions of the lengths and angles of the bonds, we first distinguish the two types of bond lengths, which we denote $A$ and $Z$, and the two types of angles, which we denote by $\alpha$ and $\zeta$, as illustrated in Figure 1. The $A$ bonds are along the armchair direction (equivalently, perpendicular to the zigzag direction). The $Z$ bonds alternate symmetrically along the zigzag direction and both exhibit identical deformations at 0 K when a uniaxial stress is applied along a high symmetry direction, such as the zigzag or the armchair direction. The angles $\alpha$ and $\zeta$ represent the bond angles formed between two $Z$ bonds and between an $A$ and a $Z$ bond, respectively. They respond always differently under an applied stress, due to the geometry of the system and the constraint $\alpha+2\zeta=2\pi$.
When a load is applied at zero temperature, there is no variability in any type of bond length and angle due to the static nature of the strained sheet at 0 K. Approximate expressions for the strain dependence of bond lengths $A$ and $Z$ and angles $\alpha$ and $\zeta$ were determined in Ref. ref-Kalosakas-Mat-2022 . Indicating by the indices $a$ or $z$ a load applied along the armchair or zigzag direction, respectively, these expressions read
@@
In Equations (8), (9), (12) and (13), the constants have units Å, the coefficients of $\epsilon_{0}$ have units Å/(%strain), and the coefficients of $\epsilon_{0}^{2}$ have units Å/(%strain)2. Similarly, in Equations (10), (11), (14) and (15) the constants have units of deg, the coefficients of $\epsilon_{0}$ have units deg/(%strain), and the coefficients of $\epsilon_{0}^{2}$ have units of deg/(%strain)2.
@@
The distributions of all the bond lengths and angles in the bulk graphene sheet at $T=0$ K are given by double singular peaks at the locations provided by the above pairs of relations for the bond lengths and angles, respectively, for different values of the applied uniaxial strain.
@@
In order to reveal the influence of temperature on the bond lengths and angles distributions, we register all the bond length and angle values during the system’s evolution in our measurement window and obtain normalized distributions for different amounts of stress/stain at various temperatures. In particular, we create a distribution for each realization by allocating all the measured bond lengths (angles) into fine-grained bins of width $3.5\times 10^{-3}$ ($0.004\,\frac{180^{\circ}}{\pi}$). We normalize the resulting distributions and then average them over the different realizations to obtain the final distribution for each case. It is worth noting that the size of the error-bars, indicating one standard deviation of this averaging computation over the different realizations, are negligible, and hence not included in the plots of the distributions presented below. We emphasize again that we consider the central region of the sheet for collecting our data, as mentioned at the end of Section 3.1.
@@
In Figures 7(a) - (d) we show the normalized bond length distributions for increasing values of stress applied along the armchair direction. In Figures 7(e) - (h), the applied load is along the zigzag direction and the stress increases from (e) to (h) too. When there is no stress, $\sigma=0$, at finite temperatures the distributions (not shown in this figure) are simply normal distributions with the variance linearly increasing with temperature (see Section 3.3.1 below). Increased temperature leads to larger fluctuations in the lattice, resulting in a wider spread of the observed bond lengths. For the smaller values of stress presented in Figures 7(a) and (e), there is a slight skewing of the distributions, but as the stress is increased the single peak splits into two peaks which are gradually separated more and more as can be clearly seen at the lower temperatures. However, the increase of temperature leads to the merging of these two peaks due to their broadening. The centers of the peaks correspond to the zero-temperature values of the two types of bond lengths for each different direction of the applied stress as given in Equations (8) - (9) and (12) - (13).
@@
Figure 7: Normalized bond length distributions in graphene, for increasing applied stress (left to right), along the armchair (top row) and the zigzag (bottom row) direction, at different temperatures $T$ as indicated in the legend. The stresses $\sigma$ in the armchair direction are (a) 0.569 eV/Å2, (b) 0.895 eV/Å2, (c) 1.22 eV/Å2, and (d) 1.55 eV/Å2, while in the zigzag direction are (e) 0.563 eV/Å2, (f) 1.13 eV/Å2, (g) 1.69 eV/Å2, and (h) 2.16 eV/Å2.
@@
Since there are twice as many $Z$ type bond lengths as $A$ types, the highest peak in each distribution in Figure 7 is mostly encompassing the lengths of the $Z$ type bonds. Thus, we can see that for stress applied along the armchair direction [Figures 7(a) - (d)], it is the $A$ bonds which achieve greater lengths (the lower peak, further to the right) while the $Z$ bonds exhibit a smaller extension. In contrast, for stress applied along the zigzag direction [Figures 7(e) - (h)], the $Z$ bonds achieve greater lengths (the taller peak is to the right in the distributions) while the centers of the smaller peaks remain near $r_{0}=1.42$, in accordance with Equation (12) and the corresponding broadening due to thermal effects. The fact that all bonds stretch for stress applied along the armchair direction, but only the $Z$ type bonds (two-thirds of all the considered bonds) stretch for a load along the other direction ref-Kalosakas-Mat-2022 , explains why the gap between the two peaks is more pronounced for stress applied in the zigzag direction.
@@
In Figure 8 we present similar results as in Figure 7, but for the distribution of bond angles. Again we note the peak splitting due to increased stress, while increasing the temperature leads to the broadening and merging of these peaks. At zero strain normal distributions centered about the equilibrium angle of $\phi_{0}=120^{\circ}$ are obtained, with a variance increasing with temperature (see Section 3.3.1). The highest peak in the bond angle distributions corresponds to the $\zeta$ angle, since there are twice as many $\zeta$ angles as $\alpha$ angles. For stress along the armchair direction [Figures 8(a) - (d)], the $\alpha$ angles decrease, while the $\zeta$ angles increase. Reverse is the situation when the stress is applied along the zigzag direction [Figures 8(e) - (h)]. We again note that the peaks of the distributions are centered about the zero temperature $\alpha$ and $\zeta$ values, as given in Equations (10) - (11) and (14) - (15), depending on the direction of the applied stress.
@@
Figure 8: Normalized bond angle distributions in planar graphene, for increasing stress (left to right), applied along the armchair (top row) and in the zigzag (bottom row) direction, at various temperatures $T$ as indicated in the legend. The values of stress $\sigma$ along the armchair direction are (a) 0.244 eV/Å2, (b) 0.651 eV/Å2, (c) 1.06 eV/Å2, and (d) 1.55 eV/Å2, while in the zigzag direction are (e) 0.282 eV/Å2, (f) 0.939 eV/Å2, (g) 1.60 eV/Å2, and (h) 2.16 eV/Å2.
3.3.1 Analytical expressions of bond length and bond angle distributions
@@
We now present approximate analytical expressions for the bond length and angle distributions, as those shown in Figures 7 and 8, in order to describe the dependence of graphene’s structural properties on stress and temperature. Based on the results discussed in the previous Subsection, we note that the obtained distributions appear to approximately be given through the combination of two normal distributions, where the means of these normal distributions correspond to the two types of bond lengths, or angles, found for each stress at zero temperature. The variance of these normal distributions is induced by thermal fluctuations, while the difference in peak heights is related to the fact that there exist double as many of one type of bond length (or angle) as the other.
@@
As there exist approximate expressions available for the equilibrium bond lengths and bond angles as a function of the applied strain at $T=0$ K [see Equations (8) - (15)], it remains to determine the explicit dependence of variance on temperature. This will be obtained by numerically evaluating the effects of temperature on the normal distributions of the bond lengths and bond angles, at the unstrained graphene sheet. We can compare the results of these calculations with analytical estimates of the variance through the Boltzmann distribution, using a second-order approximation on the relevant potential energy terms describing bond stretching and angle bending.
@@
Performing a Gaussian curve fitting procedure to the numerically obtained distributions of the bond lengths and bond angles at zero applied stress for various temperatures, shown in Figures 9(a) and (b), respectively, we compute the corresponding variances and mean values. The dependence of these variances on temperature are presented by filled circles in Figures 9(c) and (d) for the bond length and angle distributions, respectively. Solid lines in the latter plots denote a linear fitting of the data. It is worth noting that the mean of the bond length distribution slightly increases with temperature too. This is due to the thermal expansion of the 2D lattice. However, incorporating this small variation of the mean value with temperature does not practically affect the results discussed below. The mean of the bond angle distribution does not change with temperature, as expected.
@@
Figure 9: Normalized distributions of (a) bond lengths and (b) bond angles in bulk graphene for $\sigma=0$, at different temperatures $T$ as indicated in the legend. (c) data points show the dependence on $T$ of the computed variance $\Sigma_{M}^{2}$ of the bond length distributions presented in (a). Similarly, data points in (d) denote the variance $\Sigma_{B}^{2}$ of the bond angle distributions shown in (b) versus temperature. Solid lines in (c) and (d) indicate linear fitting of the corresponding data, see Equations (16) and (17), while dotted lines denote the analytical approximating expression of Equations (19) and (20).
@@
The numerically determined linear fitting of the variance $\Sigma^{2}_{M}$ of bond length distributions with the temperature $T$ is
@@
with $C_{M}=1.66\times 10^{-6}$2/K [solid line in Figure 9(c)]. Similarly, the numerically found variances $\Sigma^{2}_{B}$ for the bond angle distributions are well described by
@@
The linear dependence of these variances on temperature can be derived by the Boltzmann distribution when a quadratic approximation of the corresponding potential energy is considered. In particular, by the second derivative of the Morse potential of Equation (1),
@@
the second order approximation of the bond stretching energy term about $r=r_{0}$ reads $V_{M}^{lin}=a^{2}D(r-r_{0})^{2}$. Using this approximation, the corresponding Boltzmann distribution $\exp\left(\frac{-a^{2}D(r-r_{0})^{2}}{k_{B}T}\right)$ results in a normal bond length distribution of the form $\exp\left(\frac{-(r-r_{0})^{2}}{2\Sigma_{M}^{2}}\right)$ centered about the mean $r_{0}$ with a variance
@@
which gives $\Sigma^{2}_{M}(T)=1.97\times 10^{-6}\,T$, where the coefficient of $T$ has units of Å2/K.
Following a similar approach to estimate the variance of the bond angle distributions for different temperatures, we consider the second-order approximation of the potential $V_{B}$, Equation (2), about $\phi=\phi_{0}=120^{\circ}$, given by $V_{B}^{lin}=\frac{d}{2}\frac{\pi^{2}}{180^{2}}(\phi-\phi_{0})^{2}$ (when angles are measured in degrees). Note that due to the constrains in the sums of the bond angles around a particular atom and within hexagonal rings, just one angle can not be varied alone, but at least three other angles also change. Thus, we incorporate a factor 4 into the obtained energy approximation in the Boltzmann distribution, and eventually find that the the variance (in degrees), about the mean $\phi_{0}$, of the bond angle distribution is
@@
which results in $\Sigma^{2}_{B}(T)=1.01\times 10^{-2}\,T$, where the coefficient of $T$ has units of deg2/K.
Dotted lines in Figures 9(c) and (d) correspond to the analytical expressions of Equations (19) and (20), respectively. We can see from Figure 9(c) that the analytically predicted slope of Equation (19) is somehow larger than the corresponding numerical value (the relative difference is less than 20%). Concerning the variance of the bond angle distributions, Figure 9(d) shows an excellent agreement between the analytically and numerically obtained slopes, exhibiting a relative difference of less than 1%. The reason for the quantitative disparity between the analytical prediction and numerical determination of the slope in the linear temperature dependence of the variance of the bond length distributions, but not for the angle distributions, may be due to the fact that the second order approximation of the angle bending potential $V_{B}(\phi)$ of Equation (2) is valid for a wide range of angles (see Figure 2 in Ref. ref-Kalosakas-JAP-2013 ). On the other hand, due to the highly anharmonic nature of the Morse potential $V_{M}(r)$, Equation (1), in the same energy scales (see Figure 1 in Ref. ref-Kalosakas-JAP-2013 ), the second order approximation about $r_{0}$ is only valid very close to $r_{0}$. Thus for small temperatures, for which the achieved bond lengths are all nearby $r_{0}$, the analytical and numerical predictions in Figure 9(c) are quite close. Nevertheless, as the temperature $T$ increases, much larger bond lengths come into play and hence the analytical prediction of the variances significantly deviates from the numerical results.
Combining now the numerically determined variances for different temperatures and the known bond length and angle mean values as a function of the applied stress/strain, analytical approximate expressions for the bond length and angle distributions can be derived. Regarding the bond length distributions, an additional issue should be taken into account when the numerically determined variances from Equation (16) will be used. In particular, the relation between the variance and the temperature should be scaled according to the behavior of $V^{\prime\prime}_{M}$, Equation (18), at the mean of the corresponding peak of the distribution, since bond lengths even further away from $r_{0}$ are encountered once stress is applied to the system. Given that analytically the variance equates to $k_{B}T/V_{M}^{\prime\prime}(r_{0})$ close to $r=r_{0}$, we multiply the numerically determined variance from Equation (16) by the scaling function
@@
where $r$ is the known mean of the peak of interest in the distribution.
@@
$P_{M}=\frac{1}{\sqrt{2\pi\,C_{M}F\left(A\right)\,T}}\exp\left(-\frac{(r-A)^{2}}{2\,C_{M}F\left(A\right)\,T}\right)+\frac{2}{\sqrt{2\pi\,C_{M}F\left(Z\right)\,T}}\exp\left(-\frac{(r-Z)^{2}}{2\,C_{M}F\left(Z\right)\,T}\right),$ (22)
@@
where $A$ and $Z$ are functions of the applied stress/strain, determined in Equations (8) - (9) and Equations (12) - (13), respectively, $C_{M}$ is given in Equation (16), and $F(r)$ is provided by Equation (21). Note that the quantities $A$ and $Z$ are provided by the corresponding zero temperature relations in Equations (8), (9), (12), and (13) as a function of strain $\epsilon$. If they are needed as a function of stress $\sigma$, the stress-strain relation of Equation (7) should be used to change the variable of the applied load.
@@
For the bond angle distribution, the subtlety mentioned above concerning the scaling function is not needed since the second derivative of the angle bending potential, Equation (2), is everywhere the same regardless of the angle value at the peak of the distribution. Therefore, the angle bending distributions can be approximated by the expression
$P_{B}=\frac{1}{\sqrt{2\pi\,C_{B}\,T}}\left[\exp\left(\frac{-(\phi-\alpha)^{2}}{2\,C_{B}\,T}\right)+2\,\exp\left(\frac{-(\phi-\zeta)^{2}}{2\,C_{B}\,T}\right)\right],$ (23)
where $\alpha$ and $\zeta$ are determined by the applied stress/strain from Equations (10), (11), (14), (15), and Equation (7) if necessary, while $C_{B}$ is given in (17).
@@
Points in Figure 10 present the numerically computed bond length distributions at various applied stresses and temperatures, while the solid lines correspond to the normalized curve $P_{M}$ from Equation (22). Figure 11 contains similar results, but for the bond angle distributions. From the plots of Figure 11 we see that the analytical expression $P_{B}$, Equation (23), describes the data very well, apart from small discrepancies at the heights of the taller peak at the lower $T$ depicted and for the smaller values of stress. On the other hand, the analytical expression $P_{M}$ from Equation (22) in Figure 10 deviates from the numerically obtained distributions for the longer-bond peak at the larger values of applied stresses [Figures 10(d), (g) and (h)]. Nonetheless, both expressions of Equations (22) and (23) provide a useful analytical description of the underlying structural properties of the strained graphene at finite temperatures.

### v4

@@ 3.3 Bond length and bond angle distributions
In order to analyze the effects of temperature and stress on the distributions of the lengths and angles of the bonds, we first distinguish the two types of bond lengths, denoted by $A$ and $Z$, and the two types of angles, indicated by $\alpha$ and $\zeta$, as illustrated in Figure 1. The $A$ bonds are along the armchair direction. The $Z$ bonds alternate symmetrically along the zigzag direction and both exhibit identical deformations at 0 K when a uniaxial stress is applied along the high symmetry zigzag or armchair directions. The angles $\alpha$ and $\zeta$ represent the bond angles formed between two consecutive $Z$ bonds and between an $A$ and a $Z$ bond, respectively. They respond always oppositively under an applied stress, due to the geometry of the system and the constraint $\alpha+2\zeta=2\pi$.
When a load is applied at zero temperature, due to the absence of fluctuations and the static nature of the strained sheet, there is no variability in the two types of bond lengths and angles and their distribution is delta-peaked. Approximate expressions for the strain dependence of bond lengths $A$ and $Z$ and angles $\alpha$ and $\zeta$ were provided in Reference ref-Kalosakas-Mat-2022 . Indicating by the indices $a$ or $z$ a load applied along the armchair or zigzag direction, respectively, these expressions read
@@
In Equations (8), (9), (12) and (13), the bond lengths $A$ or $Z$ are given in Å and the zero temperature strain $\epsilon_{0}$ is expressed as % strain. Similarly, in Equations (10), (11), (14) and (15) the bond angles $\alpha$ and $\zeta$ are provided in degrees and $\epsilon_{0}$ should be given again in percentage strain. The distributions of bond lengths (angles) in bulk graphene at $T=0$ K are given by double singular peaks at the locations provided by the above pairs of relations for the bond lengths (angles), depending on the direction of the loading, for different values of the applied uniaxial strain.
@@
In order to reveal the influence of temperature on the bond lengths and angles distributions, we register all the fluctuating bond length and angle values during the system’s evolution in our measurement window and obtain normalized distributions for different amounts of stress/stain at various temperatures. In particular, we create a distribution for each realization by allocating all the measured bond lengths (angles) into fine-grained bins of width $3.5\times 10^{-3}$ Å ($0.004\,\frac{180^{\circ}}{\pi}$). The resulting distributions are normalized and then averaged over the different realizations in order to obtain the final distribution for each case. It is worth noting that the size of the error-bars, indicating one standard deviation of this averaging computation over the different realizations, are negligible, and hence not included in the plots of the distributions presented below. We emphasize again that we consider the central region of the sheet for collecting our data, as mentioned at the end of Section 3.1.
@@
In Figures 7(a) - (d) we show the normalized bond length distributions for increasing values of stress applied along the armchair direction. In Figures 7(e) - (h), the applied load is along the zigzag direction and the stress increases from (e) to (h) too. When there is no loading, $\sigma=0$, at finite temperatures the distributions are simply normal distributions with the variance linearly increasing with temperature (see Figure 9(a) below). Increased temperature leads to larger fluctuations in the lattice, resulting in a wider spread of the observed bond lengths. In the presence of uniaxial loading, for the smaller values of stress presented in Figures 7(a) and (e), there is a slight skewing of the distributions. As the stress is increased the single peak splits into two peaks which are gradually separated more and more as can be clearly seen from the plots corresponding to lower temperatures, due to the increased separation of the $A$ and $Z$ bond length values (see Figures 2 and 3 of Reference ref-Kalosakas-Mat-2022 ). However, the increase of temperature leads to the merging of these two peaks due to their broadening. The centers of the peaks correspond to the zero-temperature values of the two types of bond lengths for each different direction of the applied stress as given in Equations (8) - (9) and (12) - (13).
@@
Figure 7: Normalized bond length distributions in graphene, for increasing applied stress (left to right), along the armchair (top row) and the zigzag (bottom row) direction, at different temperatures $T$ as indicated in the legend. The curves are guides to the eyes. The vertical lines indicate the values of the $A$ and $Z$ bond lengths at zero temperature, from Equations (8), (9), (12), and (13). The stresses $\sigma$ in the armchair direction are (a) 0.569 eV/Å2, (b) 0.895 eV/Å2, (c) 1.22 eV/Å2, and (d) 1.55 eV/Å2, while in the zigzag direction are (e) 0.563 eV/Å2, (f) 1.13 eV/Å2, (g) 1.69 eV/Å2, and (h) 2.16 eV/Å2.
@@
Since there are twice as many $Z$ type bond lengths as $A$ types, the highest peak in each distribution in Figure 7 is mostly encompassing the lengths of the $Z$ type bonds. Thus, we can see that for stress applied along the armchair direction [Figures 7(a) - (d)], it is the $A$ bonds which achieve greater lengths (the lower peak, further to the right) while the $Z$ bonds exhibit a smaller extension. In contrast, for stress applied along the zigzag direction [Figures 7(e) - (h)], the $Z$ bonds achieve greater lengths (the taller peak is to the right in the distributions) while the centers of the smaller peaks remain near $r_{0}=1.42$ Å, in accordance with Equation (12) and the corresponding broadening due to thermal effects. The fact that all bonds stretch for stress applied along the armchair direction, but only the $Z$ type bonds (two-thirds of all the considered bonds) are extended for a load along the other direction ref-Kalosakas-Mat-2022 , justifies that the gap between the two peaks is more pronounced for stress applied in the zigzag direction.
@@
In Figure 8 similar results are presented as in Figure 7, but for the distribution of bond angles. At zero strain, normal distributions centered about the equilibrium angle of $\phi_{0}=120^{\circ}$ are obtained for finite temperatures, with a variance increasing with temperature (see Figure 9(b) below). Again we see the gradual peak splitting due to increased stress, while increasing the temperature leads to the broadening and merging of these peaks. The highest peak in the bond angle distributions corresponds to the $\zeta$ type angles, since there are twice as many $\zeta$ angles as $\alpha$ angles. For stress along the armchair direction [Figures 8(a) - (d)], the $\alpha$ angles decrease, while the $\zeta$ angles increase. Reverse is the situation when the stress is applied along the zigzag direction [Figures 8(e) - (h)]. Also in this case the peaks of the distributions are centered about the zero temperature $\alpha$ and $\zeta$ values, as given in Equations (10) - (11) or (14) - (15), depending on the direction of the applied stress.
@@
Figure 8: Normalized bond angle distributions in planar graphene, for increasing stress (left to right), applied along the armchair (top row) and in the zigzag (bottom row) direction, at various temperatures $T$ as indicated in the legend. The curves are guides to the eyes. The vertical lines indicate the values of the $\alpha$ and $\zeta$ bond angles at zero temperature, from Equations (10), (11), (14), and (15). The values of stress $\sigma$ along the armchair direction are (a) 0.244 eV/Å2, (b) 0.651 eV/Å2, (c) 1.06 eV/Å2, and (d) 1.55 eV/Å2, while in the zigzag direction are (e) 0.282 eV/Å2, (f) 0.939 eV/Å2, (g) 1.60 eV/Å2, and (h) 2.16 eV/Å2.
3.3.1 Analytical expressions for the bond length and bond angle distributions
@@
We now present approximate analytical expressions for the bond length and angle distributions, as those shown in Figures 7 and 8, in order to describe the dependence of graphene’s structural properties on stress and temperature. Based on the results discussed in the previous subsection, we note that the obtained distributions appear to approximately be given through the combination of two normal distributions, where the means of these normal distributions correspond to the values of the two types of bond lengths, or angles, found for each stress at zero temperature. The variance of these normal distributions is induced by thermal fluctuations, while the difference in peak heights is related to the fact that there exist double as many of one type of bond length (or angle) as the other.
@@
As there exist approximate expressions available for the equilibrium bond lengths and bond angles as a function of the applied strain at $T=0$ K [see Equations (8) - (15)], it remains to determine the explicit dependence of variance on temperature. This will be obtained by numerically evaluating the effects of temperature on the normal distributions of the bond lengths and bond angles, at the unstrained graphene sheet. The results of these calculations are compared with analytical estimates of the variance through the Boltzmann distribution, using a second-order approximation on the relevant potential energy terms describing bond stretching and angle bending.
@@
Performing a Gaussian curve fitting procedure to the numerically obtained distributions of the bond lengths and bond angles at zero applied stress for various temperatures, shown in Figures 9(a) and (b), respectively, we compute the corresponding variances and mean values. The dependence of these variances on temperature are presented by filled circles in Figures 9(c) and (d) for the bond length and angle distributions, respectively. Solid lines in the latter plots denote a linear fitting of the data. It is worth noting that the mean of the bond length distribution slightly increases with temperature too, due to the soft Morse potential describing bond stretching. However, incorporating this small variation of the mean value with temperature does not practically affect the results discussed below. The mean of the bond angle distribution does not change with temperature, as expected due to the equality of the $\alpha$ and $\zeta$ angles in the unstrained graphene and their constrained sum.
@@
Figure 9: Normalized distributions of (a) bond lengths and (b) bond angles in bulk unstrained graphene ($\sigma=0$), at different temperatures $T$, as indicated in the legend, are shown by symbols. Solid lines in (a) and (b) represent Gaussian fittings of the numerical data. Circles present the temperature dependence of (c) the variance $\Sigma_{M}^{2}$ of the Gaussian fitting of the bond length distributions presented in (a) and (d) the variance $\Sigma_{B}^{2}$ of the Gaussian fitting of the bond angle distributions shown in (b). Solid lines in (c) and (d) indicate linear fittings of the corresponding data, see Equations (16) and (17), while dotted lines denote the analytical approximating expressions of Equations (19) and (20), respectively.
@@
A linear fitting describes the dependence of the variance $\Sigma^{2}_{M}$ of bond length distributions on temperature $T$
@@
with $C_{M}=1.66\times 10^{-6}$ Å2/K [solid line in Figure 9(c)]. Similarly, the numerically found variances $\Sigma^{2}_{B}$ for the bond angle distributions are well described by
@@
The proportionality of these variances with temperature can be derived through the Boltzmann distribution when a quadratic approximation of the corresponding potential energy is considered. In particular, by the second derivative of the Morse potential of Equation (1),
@@
the second order approximation of the bond stretching energy term about the equilibrium $r=r_{0}$ reads $V_{M}^{lin}=a^{2}D(r-r_{0})^{2}$. Using this approximation, the corresponding Boltzmann distribution $\exp\left(\frac{-a^{2}D(r-r_{0})^{2}}{k_{B}T}\right)$ results in a normal bond length distribution of the form $\exp\left(\frac{-(r-r_{0})^{2}}{2\Sigma_{M}^{2}}\right)$ centered about the mean $r_{0}$ with a variance
@@
which gives $\Sigma^{2}_{M}(T)=1.97\times 10^{-6}$2/K) $\cdot\;T$ when the parameters of the Morse potential are substituted.
Following a similar approach to estimate the variance of the bond angle distributions for different temperatures, we consider the second-order approximation of the potential $V_{B}$, Equation (2), about $\phi=\phi_{0}=120^{\circ}$, given by $V_{B}^{lin}=\frac{d}{2}\frac{\pi^{2}}{180^{2}}(\phi-\phi_{0})^{2}$ (when angles are measured in degrees). Note that due to the constrains in the sums of the bond angles around a particular atom and also in the sums within hexagonal rings, just one angle can not be varied alone. When an angle slightly varies from the equilibrium value, at least three other angles should be also changed. Thus, multiplying by a factor 4 the linearized angle bending energy in the Boltzmann distribution, we eventually find the variance (in squared degrees), about the mean $\phi_{0}$, of the bond angle distribution
@@
which results in $\Sigma^{2}_{B}(T)=1.01\times 10^{-2}$ (deg2/K) $\cdot\,T$, using the value of $d$.
Dotted lines in Figures 9(c) and (d) correspond to the analytical expressions of Equations (19) and (20), respectively. We can see from Figure 9(c) that the analytically obtained slope of Equation (19) is somehow larger than the corresponding numerical value of Equation (16) (the relative difference is less than 20%). Concerning the variance of the bond angle distributions, Figure 9(d) shows an excellent agreement between the analytically and numerically obtained slopes, exhibiting a relative difference of less than 1%. One reason for the quantitative disparity between the analytical prediction and numerical determination of the slope in the linear temperature dependence of the variance of the bond length distributions, but not for the angle distributions, may be due to the fact that the second order approximation of the angle bending potential $V_{B}(\phi)$ of Equation (2) is valid for a wide range of angles (see Figure 2 in Reference ref-Kalosakas-JAP-2013 ). On the other hand, due to the highly anharmonic nature of the Morse potential $V_{M}(r)$, Equation (1), in the same energy scales (see Figure 1 in Reference ref-Kalosakas-JAP-2013 ), the second order approximation about $r_{0}$ is only valid very close to $r_{0}$.
Combining now the numerically determined variances for different temperatures and the known bond length and angle mean values as a function of the applied stress/strain, analytical approximate expressions for the bond length and angle distributions can be derived. Regarding the bond length distributions, an additional issue should be taken into account when the numerically determined variances from Equation (16) will be used. In particular, the relation between the variance and the temperature should be scaled according to the behavior of $V^{\prime\prime}_{M}$, Equation (18), at the mean of the corresponding peak of the distribution, since bond lengths even further away from $r_{0}$ are encountered once stress is applied to the system and the second derivative of the Morse potential varies significantly with $r$. Given that analytically the variance equates to $k_{B}T/V_{M}^{\prime\prime}(r_{0})$ close to $r=r_{0}$, we multiply the numerically determined variance from Equation (16) by the scaling function
@@
where $r$ is the known mean of the peak of interest in the distribution, provided by either Equations (8) - (9) or Equations (12) - (13) depending on the loading direction.
@@
$P_{M}=\frac{1}{3\sqrt{2\pi\,C_{M}F\left(A\right)\,T}}\exp\left(-\frac{(r-A)^{2}}{2\,C_{M}F\left(A\right)\,T}\right)+\frac{2}{3\sqrt{2\pi\,C_{M}F\left(Z\right)\,T}}\exp\left(-\frac{(r-Z)^{2}}{2\,C_{M}F\left(Z\right)\,T}\right),$ (22)
@@
where $A$ and $Z$ are functions of the applied stress/strain, determined in Equations (8) - (9) or Equations (12) - (13), for stress along the armchair or zigzag direction respectively, $C_{M}$ is given in Equation (16), and $F(r)$ is obtained by Equation (21). The factor of 2 in the second term is because there are double $Z$ bonds than $A$ bonds. The division by 3 is for normalizing the distribution. Note that the quantities $A$ and $Z$ are provided by the corresponding zero temperature relations in Equations (8), (9), (12), and (13) as a function of strain $\epsilon$. If they are needed as a function of stress $\sigma$, the stress-strain relation of Equation (7) should be used to change the variable of the applied load.
@@
For the bond angle distributions, the subtlety mentioned above concerning the scaling function is not needed since the second derivative of the angle bending potential, Equation (2), is everywhere the same regardless of the angle value at the peak of the distribution. Therefore, the angle bending distributions can be approximated by the expression
$P_{B}=\frac{1}{3\sqrt{2\pi\,C_{B}\,T}}\left[\exp\left(-\frac{(\phi-\alpha)^{2}}{2\,C_{B}\,T}\right)+2\,\exp\left(-\frac{(\phi-\zeta)^{2}}{2\,C_{B}\,T}\right)\right],$ (23)
where $\alpha$ and $\zeta$ are determined by the applied stress/strain from Equations (10) - (11) or (14) - (15), depending on the loading direction, and $C_{B}$ is provided by Equation (17). The factor 2 in the second term is due to there being twice as many $\zeta$ angles as $\alpha$ angles and the division by 3 normalizes the distribution. If the loading is given through the value of stress, Equation (7) can be also used.
@@
Circles in Figure 10 present the numerically computed bond length distributions at various applied stresses and temperatures, while the solid lines correspond to the curve $P_{M}$ from Equation (22). Figure 11 contains similar results, but for the bond angle distributions. These plots show that the analytical expressions presented above provide overall a reasonable description of the bond length and angle distributions in strained graphene, at various temperatures at least up to the values considered here. In Figure 10, at the larger values of applied stress and lower temperatures the analytical distribution $P_{M}$ of Equation (22) underestimates the longer-bond (second) peak of the numerically obtained distributions [Figures 10(d), (g) and (h)]. In Figure 10(h) we observe the greatest deviation of the analytical expression from the numerical data at the right-hand peak of the lowest temperature at $T=100$ K; in this case the difference is 8.7%. From the plots of Figure 11 we see that the analytical expression $P_{B}$, Equation (23), describes the data quite well, apart from small discrepancies at the heights of the taller peak at the lower temperatures depicted and for the smaller values of stress. In Figure 11(a) we observe the biggest deviation for the $T=100$ K case, where the value of the analytical expression is 4.1% below that of the numerical data. In any case, both expressions of Equations (22) and (23) provide a useful analytical description of the underlying structural properties of the strained graphene at finite temperatures.